# Touch: Enabling Accurate, Lightweight Self-Touch Sensing with Passive Magnets

### v1

@@ Abstract
This paper presents Touch, a novel magnetic sensing platform for self-touch gesture recognition. Touch features (1) a compact hardware design with low-power magnetometers and magnetic silicon, (2) a lightweight semi-supervised framework requiring minimal user data, and (3) an ambient field detection module to mitigate environmental interference. We evaluated Touch in two representative applications in user studies with 11 and 12 participants. Touch only requires three-second fine-tuning data for each gesture — new users need less than one minute before starting to use the system. Touch can distinguish eight different face-touching behaviors with an average accuracy of 93.41%, and reliably detect body-scratch behaviors with an average accuracy of 94.63%. Touch demonstrates accurate and robust sensing performance even after a month, showcasing its potential as a practical tool for hygiene monitoring and dermatological health applications.
@@ IV Hardware Design of Touch
We conducted a simulation-guided computational design study to explore the trade-off between sensor quantity and classification performance. In this framework, sensors were virtually placed along the x-axis at 0.8 cm intervals, and eight representative magnet movement directions were defined across a plane ({, , , …}). Each trajectory started and ended 5 cm from the origin and was sampled 100 times with Gaussian noise (1 cm standard deviation) to emulate natural variations. A classifier was trained with 10 samples per action and evaluated on the remaining 90.
@@ VI-B Data Collection Protocol
(3) Testing: Each participant completed five sessions. In each session, they performed all gesture categories once (face: seven regions + no touching; body: scratching + no scratching). To assess robustness under varying environmental magnetic fields, each session was conducted in one of five evenly spaced orientations around . Gesture durations were not fixed, allowing natural variation, i.e., face: 3 - 15s; body: 2 - 12s.
@@ VI-C Classification Performance
For the face-touching detection task, participants using their dominant hand achieved notably higher accuracy (96.7%) than those using their non-dominant hand (89.5%) when classifying eight categories. This discrepancy likely arises because non-dominant hand movements were less consistent with training data, reflecting limited dexterity in replicating subtle face-touching gestures. Nevertheless, all non-dominant results still exceeded 89%, confirming that Touch remains reliable even in this setting.
For the body-scratch detection task, we designed a mirrored protocol to accommodate regions that cannot be scratched symmetrically (e.g., scratching the left arm with the left hand). Here, the impact of hand dominance was less pronounced: both dominant and non-dominant hand conditions maintained accuracy above 94%. This suggests that habitual scratching behaviors are less constrained by hand dexterity than face-touching, making Touch broadly effective regardless of hand dominance.
@@ VI-C2 Ablation Study
As shown in Table II, removing MagDelta led to a moderate accuracy drop in the face-touching task (from 93.41% to 88.18%). Here, the stronger magnetic field generated by the magnetic rings still provided separable signals, even though robustness across orientations degraded.
@@ VI-C3 Remount Experiment
These results demonstrate that Touch is resilient to minor placement variations and remains reliable even when reattached by end-users, including patients with real dermatological conditions, without expert calibration.
@@ VI-D Energy Efficiency and Latency
We measured the power consumption of Touch by monitoring the voltage and current drawn from its battery under continuous operation. The sensing unit, which includes three magnetometers, the control circuit, and the BLE module, consumed on average 25 mW. With a compact 3.7 V, 100 mAh battery, this corresponds to approximately 8 hours of uninterrupted operation, demonstrating suitability for daily use.
@@ VII-A Self-touch Detection
Prior work on self-touch detection schemes has explored diverse sensing modalities for detecting face-touching events [7, 21]. COVID-away [35] employs a multi-sensor wearable that extracts over 100 features from motion and pressure signals, but it only distinguishes touching versus non-touching, without fine-grained recognition. SavingFace [31] leverages acoustic sensing through wired earbuds but is limited to coarse detection. FaceSense [14] combines thermal imaging with physiological signals (impedance, EMG), yet primarily addresses binary classification and requires bulky hardware. D-Touch [19] achieves finer granularity with a neck-mounted infrared camera, recognizing up to 17 facial activities, but at the cost of high power consumption and privacy concerns.
These approaches highlight the trade-offs of existing modalities. That is, while some achieve fine-grained classification, they often rely on intrusive, power-hungry, or privacy-sensitive sensors; others remain restricted to binary detection.
@@ VII-B Magnetic Sensing
Electromagnet Tracking. Electromagnetic tracking offers high accuracy, as alternating magnetic fields are stable and resistant to environmental noise. Finexus [4] leverages multiple electromagnetic frequencies and filters to track finger movements, while AuraRing [27] uses a coil in a ring and magnet sensors in a wristband to achieve 3-DoF position and 2-DoF orientation tracking with iterative models anda neural network. However, electromagnetic tracking has notable drawbacks compared to Touch’s battery-free solution. It requires a powered coil on the finger, thus extensively increasing energy consumption and generating substantial heat, with coils potentially reaching temperatures high enough to melt 3D-printed PLA enclosures (melting point around C). The receiver is usually bulkier compared with Touch. For example, AuraRing [27] needs a wrist-sized receiver.

### v2

@@ Abstract
This paper presents Touch, a novel magnetic sensing platform for self-touch gesture recognition. Touch features (1) a compact hardware design with low-power magnetometers and magnetic silicon, (2) a lightweight semi-supervised framework requiring minimal user data, and (3) an ambient field detection module to mitigate environmental interference. We evaluated Touch in two representative applications in user studies with 11 and 12 participants. Touch only requires three-second fine-tuning data for each gesture — new users need less than one minute before starting to use the system. Touch can distinguish eight different face-touching behaviors with an average accuracy of 93.41%, and reliably detect body-scratch behaviors with an average accuracy of 94.63%. Touch demonstrates accurate and robust sensing performance even after a month, showcasing its potential as a practical tool for hygiene monitoring and dermatological health applications. Code is available at https://wangmerlyn.github.io/muTouch/.
@@ IV Hardware Design of Touch
We conducted a simulation-guided computational design study to explore the trade-off between sensor quantity and classification performance. In this framework, sensors were virtually placed along the x-axis at 0.8 cm intervals, and eight representative magnet movement directions were defined across a 360° plane ({, 45°, 90°, …}). Each trajectory started and ended 5 cm from the origin and was sampled 100 times with Gaussian noise (1 cm standard deviation) to emulate natural variations. A classifier was trained with 10 samples per action and evaluated on the remaining 90.
@@ VI-B Data Collection Protocol
(3) Testing: Each participant completed five sessions. In each session, they performed all gesture categories once (face: seven regions + no touching; body: scratching + no scratching). To assess robustness under varying environmental magnetic fields, each session was conducted in one of five evenly spaced orientations around 360°. Gesture durations were not fixed, allowing natural variation, i.e., face: 3 - 15s; body: 2 - 12s.
@@ VI-C Classification Performance
For the face-touching detection task, participants using their dominant hand achieved higher accuracy (96.7%) than those using non-dominant hand (89.5%). This discrepancy likely arises because non-dominant hand movements were less consistent with training data, reflecting limited dexterity in replicating subtle face-touching gestures. Nevertheless, all non-dominant results still exceeded 89%, confirming that Touch remains reliable in this setting.
For body-scratch detection, we designed a mirrored protocol to accommodate regions that cannot be scratched symmetrically (e.g., scratching the left arm with the left hand). Here, the impact of hand dominance was less pronounced: both dominant and non-dominant hand conditions maintained accuracy above 94%.
@@ VI-C2 Ablation Study
As shown in Table II, removing MagDelta led to a moderate accuracy drop in face-touching (from 93.41% to 88.18%). Here, the stronger magnetic field generated by the magnetic rings still provided separable signals, even though robustness across orientations degraded.
@@ VI-C3 Remount Experiment
These results demonstrate that Touch is resilient to minor placement variations and remains reliable even when reattached by users, including patients with real dermatological conditions, without expert calibration.
@@ VI-D Energy Efficiency and Latency
We measured the power consumption of Touch by monitoring the voltage and current drawn from its battery under continuous operation. The whole sensing hardware, which includes three magnetometers, the control circuit, and the BLE module, consumed on average 25 mW. With a compact 3.7 V, 100 mAh battery, this corresponds to approximately 8 hours of uninterrupted operation, demonstrating suitability for daily use.
@@ VII-A Self-touch Detection
Prior work on self-touch detection has explored diverse sensing modalities for detecting face-touching [7, 21]. COVID-away [35] employs a multi-sensor wearable extracting over 100 features from motion and pressure signals, but only distinguishes touching from non-touching. SavingFace [31] leverages acoustic sensing via wired earbuds but is limited to coarse detection. FaceSense [14] combines thermal imaging with physiological signals (impedance, EMG), yet mainly addresses binary classification and requires bulky hardware. D-Touch [19] achieves finer granularity with a neck-mounted infrared camera, recognizing up to 17 facial activities, but at the cost of high power consumption and privacy concerns.
These approaches highlight trade-offs of existing modalities: while some achieve fine-grained classification, they rely on intrusive, power-hungry, or privacy-sensitive sensors; others remain limited to binary detection.
@@ VII-B Magnetic Sensing
Electromagnet Tracking. Electromagnetic tracking offers high accuracy, as alternating magnetic fields are stable and resistant to environmental noise. Finexus [4] leverages multiple electromagnetic frequencies and filters to track finger movements, while AuraRing [27] uses a coil in a ring and magnet sensors in a wristband to achieve 3-DoF position and 2-DoF orientation tracking with iterative models anda neural network. However, electromagnetic tracking has notable drawbacks compared to Touch’s battery-free solution. It requires a powered coil on the finger, thus extensively increasing energy consumption and generating substantial heat, with coils potentially reaching temperatures high enough to melt 3D-printed PLA enclosures (melting point around 180°C). The receiver is usually bulkier compared with Touch. For example, AuraRing [27] needs a wrist-sized receiver.